OpenStudy (anonymous):

The graph of f(x) = x2 is shifted 3 units to the right to obtain the graph of g(x). Which of the following equations best describes g(x)?

OpenStudy (anonymous):

g(x) = (x + 3)2 g(x) = x2 - 3 g(x) = x2 + 3 g(x) = (x - 3)2 Are the possible answers.

OpenStudy (anonymous):

Shifting a function f(x) to the right or left is done by subtracting or adding respectively the numbers from the variable x in the function. So when you shift f(x) = x^2 3 units right it becomes f(x - 3) = (x - 3)^2

OpenStudy (anonymous):

So it would be D? @navk

OpenStudy (aravindg):

yes.

OpenStudy (anonymous):

Yes it is D